If the product of 2 consecutive integers is 45 more than 5 times the first integer. What are the 2 integers?
Answer by mananth(16949)

one integer be x
other will be (x+1)
x(x+1)=5x+45
x^2+x=5x+45
x^2-5x+x-45=0
x^2-4x-45=0
x^2-9x+5x-45=0
x(x-9)+5(x-9)=0
(x-9)(x+5)=0
x=9 OR -5
The integers are
(9,10) , (-5,-4)